OpenStudy (anonymous):

What is the equation of a line that passes through (-6, 2) and has a slope of −12 ?

OpenStudy (kidrah69):

@texaschic101 already showed u.... in y = mx + b form the slope is in the m position so m = -12 your points (-6,2)....x = -6 and y = 2 now we already know the slope (m) and our points (x,y)....so basically we just need to find b, the y intercept. So we sub into the formula and find b y = mx + b 2 = -12(-6) + b 2 = 72 + b 2 - 72 = b -70 = b so your equation becomes : y = -12x - 70 questions ?

OpenStudy (campbell_st):

yes, why would an ambassador give the answer, rather then explaining how...?

OpenStudy (triciaal):

@campbell_st don't just read the last line. It is explained.

OpenStudy (campbell_st):

lol... I read the who thing... that's why I commented

OpenStudy (kidrah69):

Im not a person who gives direct answers. And this was on a previous post i only copied and pasted it. And gave credit to the work of whose it was

OpenStudy (kidrah69):

Thanks.